### v1

# Emotion Granularity from Text: An Aggregate-Level Indicator of Mental Health

Abstract
We are united in how emotions are central to shaping our experiences; and yet, individuals differ greatly in how we each identify, categorize, and express emotions. In psychology, variation in the ability of individuals to differentiate between emotion concepts is called emotion granularity (determined through self-reports of one’s emotions). High emotion granularity has been linked with better mental and physical health; whereas low emotion granularity has been linked with maladaptive emotion regulation strategies and poor health outcomes. In this work, we propose computational measures of emotion granularity derived from temporally-ordered speaker utterances in social media (in lieu of self-reports that suffer from various biases). We then investigate the effectiveness of such text-derived measures of emotion granularity in functioning as markers of various mental health conditions (MHCs). We establish baseline measures of emotion granularity derived from textual utterances, and show that, at an aggregate level, emotion granularities are significantly lower for people self-reporting as having an MHC than for the control population. This paves the way towards a better understanding of the MHCs, and specifically the role emotions play in our well-being.
Emotion Granularity from Text:
An Aggregate-Level Indicator of Mental Health
[…]
1 Introduction
Emotions play a central role in how we construct meaning and communicate with those around us. Yet, individuals vary in their understanding and experience of emotions, or “emotional expertise” Hoemann et al. (2021b). Some people are able to recognize, identify, and describe what they feel using precise, context-specific terms like guilt, anger, frustration, or helplessness; others tend to use more broad terms to convey a general sense of feeling bad or feeling low.
Emotion granularity, aka emotion differentiation, is defined by psychologists as the ability of an individual to experience and categorize emotions in very specific terms Barrett et al. (2001). Highly granular individuals have a broad range of highly situated and differentiated emotion concepts, and can reliably describe these concepts using language — for example, distinguishing between when they are feeling angry vs. when they are feeling sad, or when they are feeling elated from when they are feeling content.
In psychology and the affective sciences, emotion granularity is often measured using repeated measurements, where individuals are asked to rate the intensity of experiencing certain emotions multiple times over a period of days (e.g., 2–3 times each day for 5 days). An individual’s emotional granularity is then operationalized as the extent to which multiple emotions are co-endorsed over time, i.e, how similarly the emotions are rated across all measurements. Individuals who tend to frequently rate multiple emotions at the same intensity levels are defined as low in granularity — the frequent co-endorsement across time indicates that they are failing to differentiate between these emotions in their reports. In contrast, individuals high in emotion granularity co-endorse multiple emotions less frequently over time. The level of co-endorsement is operationalized in psychology using the intraclass correlation coefficient (ICC) (Shrout and Fleiss, 1979), which measures the extent to which the emotions co-vary in reports at the aggregate level (Tugade et al., 2004; Hoemann et al., 2021a; Lee et al., 2017; Reitsema et al., 2022). Typically, granularity is measured across emotions with the same valence; one can therefore have a measure of negative emotion granularity, measured as the granularity of negative emotions (such as anger, sadness, and fear) and positive emotion granularity, measured as the granularity of positive emotions (such as joy, excitement, and satisfaction). Some works also look at the co-endorsement of emotions that express opposite valence, such as joy and sadness (Lindquist and Barrett, 2008).
Evidence collected in the last two decades provides consistent support for a link between emotional granularity and mental health Erbas et al. (2014, 2018); Starr et al. (2017); Seah et al. (2020), physical health Hoemann et al. (2021b); Bonar et al. (2023), and adaptive health behavior Dixon-Gordon et al. (2014); Kashdan et al. (2015).
Traditionally, past work in in psychology has measured emotion granularity using self-reports of emotions felt; however, there are several limitations of this approach. The number of participants and the duration over which the data can be collected are limited. Further, several biases may be present in the collected reports such as social desirability (e.g., tendency to report more desirable traits and under-report less desirable traits) (Kragel et al., 2022; Latkin et al., 2017a).
In this work, we, for the first time, propose a way to compute emotion granularity from the textual utterances of an individual. Our method uses the temporal sequence of the utterances to first construct emotion arcs along multiple emotions, and computes granularity as the correlation of these emotion arcs. We hypothesize that this measure is indicative of the individual consistently expressing the same set of emotions together over a period of time, and can therefore act as a proxy measure of emotional granularity. We generate emotion arcs for eight emotions: anger, anticipation, disgust, fear, joy, sadness, surprise, and trust using the NRC Emotion Intensity Lexicon Mohammad (2018). Similar to work in psychology, we group anger, sadness, fear, and disgust into one group (the negative valence group); joy and trust into a second group (the positive valence group); and surprise and anticipation into a third group (variable valence group). [^1] Different instances of surprise can correspond to very different amounts of valence—some positive and some negative. Same for anticipation. We compute aggregate measures of negative, positive, and variable emotional granularity; these are further aggregated to obtain an overall measure of granularity between all within-group emotion pairs.
We then study the relationship between aggregate, population-level measures of emotion granularity in text for eight Mental Health Conditions (MHCs), namely attention-deficit hyperactivity disorder (ADHD), anxiety, bipolar disorder, depression, major depressive disorder (MDD), obsessive-compulsive disorder (OCD), postpartum depression (PPD), and post-traumatic stress disorder (PTSD), and compare them to a control group. We use two social media datasets (Suhavi et al., 2022; Losada et al., 2017, 2018) where users have self-disclosed their mental health diagnosis. We compute emotion granularity metrics for each of these groups to answer the following questions:
  1. 1.
    Do measures of emotional granularity differ between the MHCs and the control group?
  2. 2.
    Which measures of emotion granularity are the most effective at differentiating between the MHCs and the control group?
  3. 3.
    Which emotion pairs lead to the greatest difference in granularity between an MHC and the control group?
Exploring this line of questions helps us better understand how emotion granularity presents itself in text, whether emotion granularity from text can be a useful tool to study MHCs, and how an MHC impacts our perception of emotions (and perhaps even, how the perception of emotions impacts our mental health). While this work is the first in exploring emotion granularity from text, and presents early results in how they correlate with MHCs, we hope it opens up an avenue of future work to substantively explore these and other questions.
We establish baseline measures of emotion granularity from text, and show that these measures function as reliable indicators, at the aggregate-level, for the presence of many of the mental health conditions we study. Our work makes an important contribution to the growing wealth of research on textual measures of emotional expression as biosocial markers of MHCs, and has a broader utility in functioning as an additional indicator of the mental well-being of populations. [^2] The term biosocial marker (Lena, 2021) was coined to indicate the crucial role social factors (e.g., socioeconomic status, years of education, bilingualism, etc.) have on quantitative features associated with medical conditions (biomarkers). We make use of publicly available datasets and all our code will be made available through the project webpage. [^3] https://github.com/Priya22/emotion-granularity
2 Related Work
2.1 Emotions and Mental Health
Measures of emotional experience and their patterns of change over time have been extensively studied as markers of mental and physical well-being Lewis et al. (2010). The Emotion Dynamics framework in psychology quantifies the patterns with which emotions change over time, allowing researchers to better understand emotional experiences and individual variation Kuppens and Verduyn (2017). The framework includes several measures such as the duration, intensity, variability, and granularity of one’s emotional experiences. Numerous studies in psychology have shown emotion dynamics correlate with overall well-being, mental health, and psychopathology (the scientific study of mental illness or disorders) (Kuppens and Verduyn, 2017; Houben et al., 2015; Silk et al., 2011; Sperry et al., 2020).
Emotion granularity in particular is positively associated with better adaptive behaviour in adverse conditions — accurately labeling our emotions can inform us of the right coping strategies to use in different contexts. Individuals with higher emotion granularity tend to use a broader range of strategies to deal with negative emotions, and are more successful at doing so Barrett et al. (2001). Several studies have shown that emotion granularity is lower in individuals with mental health conditions like bipolar disorder Suvak et al. (2011); Dixon-Gordon et al. (2014), manic depressive disorder Demiralp et al. (2012), schizophrenia Kring et al. (2003), autism spectrum disorder Erbas et al. (2013), and affective disorders like anxiety Seah et al. (2020) and depression Starr et al. (2017); Willroth et al. (2020). Lower granularity is also associated with increased tendencies to engage in maladaptive behaviour, such as alcohol consumption Kashdan et al. (2015); Emery et al. (2014) and aggression Pond Jr et al. (2012).
Researchers in affective science typically measure emotional granularity through experience sampling methodologies (ESMs), or ecological momentary assessments (EMAs), where individuals (participants) are repeatedly asked to report on their emotional states on several occasions throughout the day, for several days. Granularity is then quantified by tracking the endorsement of emotions over time. For example, participants may be asked to endorse a series of ten emotion words (e.g., "Anger," "Fear," "Happy", etc.) on a Likert scale many times across several sampling instances. Emotional granularity would then be computed as the intraclass correlation (ICC) of ratings across sampling instances. A high ICC would suggest that a participant experiences all of the emotions in a similar way across trials (treating them as synonyms for more general states like "unpleasantness" or "pleasantness"), whereas a low ICC would suggest that a participant experienced emotions in a granular and context-specific way.
While emotion granularity is generally measured between emotion categories that are close to each other in the affective space (i.e, express similar valence), the concept of dialecticism refers to the co-incidental experience of both negative and positive emotions Lindquist and Barrett (2008). Dialecticism can therefore be operationalized as the co-endorsement of emotion pairs that express positive and negative valence.
2.2 Language and Mental Health
Given the limitations of self-report surveys (e.g., limited data coverage and time spans, biases, etc. (Kragel et al., 2022)), another approach to measure well-being indicators is through one’s language usage. Some well-known linguistic indicators of mental health include the proportion of pronouns used for those with depression (Koops et al., 2023), syntax reduction for anorexia nervosa (Cuteri et al., 2022), certain lexical and syntactic features for mild cognitive impairment and dementia (Calzà et al., 2021; Gagliardi and Tamburini, 2021), and semantic connectedness for schizophrenia (Corcoran et al., 2020).
Recently, another linguistic feature that researchers leveraged for insights into overall well-being, are the emotions expressed in language. Largely, only sentiment has been explored and mainly from social media data (a rich source of language data). For example, more negative sentiment was expressed in text by individuals with depression (De Choudhury et al., 2013; Seabrook et al., 2018; De Choudhury et al., 2021). Other work has found that suicide watch, anxiety, and self-harm subreddits had markedly lower negative sentiment compared to other mental health subreddits such as Autism and Asperger’s (Gkotsis et al., 2016). Teodorescu et al. (2023) found that measures of emotion dynamics from text correlate with various mental health diagnoses. These works overall show that the average emotion expressed in text and also the characteristics of emotion change over time (e.g., variability) are meaningful indicators of well-being. In this work, we explore if measures of emotion granularity from text are likewise meaningful indicators of mental health.
3 Datasets
We use the Twitter-STMHD dataset Suhavi et al. (2022) for our experiments and also verify our results with a smaller Reddit eRisk Losada et al. (2017, 2018) dataset. We describe both of them below.
Twitter-STMHD dataset: Suhavi et al. (2022) identified tweeters who self-disclosed as having an MHC diagnosis using carefully constructed regular expression patterns and manual verification. We summarize key details on the dataset creation process in Appendix A. The control group consists of users identified from a random sample of tweets (who posted during approximately the same time period as the MHC tweets). These tweeters did not post any tweets meeting the MHC regex described above. Additionally, users who had any posts about mental health discourse were removed from the control group. Note that this process does not guarantee that users in the control group did not have an MHC diagnosis, but rather the group as a whole may have very few MHC tweeters. The number of users in the control group was selected to match the size of the depression dataset, which had the largest number of users.
For the final set of users, four years of tweets were collected for each user: two years before self-reporting a mental health diagnosis and two years after. For the control group, tweets were randomly sampled from between January 2017 and May 2021 (same date range as the other MHC classes).
Users with less than 50 tweets collected were removed so as to allow for more generalizable conclusions to be drawn. Similarly, users with more than 5000 followers were removed to not include celebrities or other organizations that use Twitter to discuss well-being.
Reddit eRisk dataset: To further add to our findings, we also included the eRisk 2018 dataset Losada et al. (2017, 2018) in our experiments. It consists of users who self-disclosed as having depression on Reddit (expressions were manually checked), and a control group (individuals were randomly sampled). The dataset includes several hundred posts per user, over approximately a 500-day period. We combined users and their instances from both the training set (which is from the eRisk 2017 task Losada et al. (2017)) and the test set (from the eRisk 2018 task Losada et al. (2018)).
3.1 Preprocessing
We further preprocessed both the Twitter-STMHD dataset and the eRisk dataset for our experiments (Section 4), as we are specifically interested in the relationship between emotion granularity and each disorder. Several users self-reported as being diagnosed with more than one disorder, referred to as comorbidity. We found a high comorbidity rate between users who self-reported as having anxiety and depression, as is also supported in the literature (Pollack, 2005; Gorman, 1996; Hirschfeld, 2001; Cummings et al., 2014). Since we wanted to focus on each MHC separately (and not on co-morbidity) we only considered users who self-reported as having one MHC. We also performed the following preprocessing steps:
  • We only considered posts in English, removing posts in other languages.
  • We filtered out posts that contained URLs (as the text in such posts is often not self-contained; and one might need to view the page pointed to by the url to get a better sense of the post text).
  • We removed retweets (identified through tweets containing ‘RT’, ‘rt’). This is to focus exclusively on texts written by the user.
  • To ensure that we did not include users that post very infrequently or very frequently we excluded those who were outliers based on the number of posts per individual. We determine outliers as those who either had less than 100 posts (as was similarly done in Vishnubhotla and Mohammad (2022)) or those who had posted more than 1.5 times the interquartile range above quartile three (75th percentile) of the control group. [^4] The interquartile range is from the 25th to 75th percentile.
  • We removed punctuation and stop words.
Table 3.1 shows key details of the filtered Twitter-STMHD and Reddit eRisk datasets.
Dataset, Group #people Av. #posts Av. #tokens
per post
Twitter
    MHC 19,324 2,590.48 17.59
        ADHD 6,356 2,497.43 17.46
        Anxiety 3,036 2,921.05 17.46
        Bipolar 1,061 2,820.17 17.32
        Depression 4,855 2,526.62 16.75
        MDD 219 2,640.69 16.40
        OCD 1,009 2,388.73 18.38
        PPD 179 2,581.19 19.18
        PTSD 2,609 2,533.85 19.41
    Control 6,001 2,420.50 16.16
Reddit
    Depression 112 556.57 47.22
    Control 907 665.00 41.09
Table 1: The number of users in each mental health condition, the average number of posts per user, and the average number of tokens per post in the preprocessed version of the Twitter-STMHD and Reddit eRisk datasets. We use this data for our experiments.
4 Emotion Granularity from Text
The core metric that we want to capture from the text utterances of an individual is emotion granularity—what psychologists term the “co-endorsement" of pairs of emotions. Analogous to their operationalization of granularity in terms of the Intra-Class Correlation (ICC) of repeated emotion intensity measurements along emotion adjectives, we use textual utterances to derive a temporal sequence of emotion states, referred to as an emotion arc for the speaker (section 4.1), and operationalize granularity as the correlations of these arcs. We construct emotion arcs for multiple emotions, for each user in the MHC groups and the control group.
Emotion Dimensions:
A key requirement of our computational method is that we must be able to quantify the emotional score of a text along a selected emotion dimension. We are therefore limited by the resources and models available to compute such a score for an emotion dimension.
Here, keeping in mind the necessity of including multiple emotion dimensions that are similarly-valenced, we work with the eight emotions represented in the NRC Emotion Intensity Lexicon: anger, anticipation, disgust, fear, joy, sadness, surprise, and trust Mohammad (2018).
We partition these emotions into three groups based on the valence association: joy and trust are in the positive valence group; anger, sadness, fear, and disgust are in the negative valence group; and anticipation and surprise are in the variable valence group. The distinction for surprise and anticipation is necessary because specific instances of these emotions can have a positive or a negative connotation (e.g., a good or a bad surprise).
4.1 Constructing Emotion Arcs
We order the utterances for each user based on timestamp information in the metadata. We construct emotion arcs for the temporal sequence of utterances of each user, along each of the eight emotions, in two ways pertaining to different window sizes. This is to make sure that the results are largely robust even when varying the window size to some extent.
Utterance-level Window:
Emotion scores (for each emotion category) are computed for each utterance (i.e, tweet or Reddit post). Here, an utterance is assumed to represent the speaker’s emotion state at a particular point in time (analogous to sampling instances). The sequence of utterance emotion scores for a user forms their temporal emotion arc. [^5] The frequency and time of posting often differs between users, but we ignore that for now.
Word-Count based Window:
Here, the emotion score at a point in time is computed for a window of words (say, 100 words) that are uttered around that point, and the window is moved forward by a fixed step size (say, 1 word at a time) to obtain the emotion score for the next time step. In prior work on constructing emotion arcs from temporally-ordered text, such sliding windows are usually employed to ensure smoother arcs that more accurately capture the flow of emotions over time.
Teodorescu and Mohammad (2023) conducted an extensive quantitative evaluation of several hyperparameters involved in emotion arc construction on datasets from diverse domains (tweets, movie reviews, diary posts) annotated with emotion scores. We follow many of their recommendation to construct emotion arcs for the utterances of each of our users.
The texts are tokenized using the twokenizer [^6] https://github.com/myleott/ark-twokenize-py package to obtain a similarly-ordered sequence of words. Emotion scores are computed with window sizes of 100 words and 500 words each, and the window is moved forward by one word at each timepoint to obtain a series of overlapping emotion scores. The final emotion score is for the window that subsumes the final word of the individual’s utterance sequence.
Emotion scoring method
Keeping in mind the necessity of an interpretable method of emotion scoring, we use word–emotion lexicons to compute the emotion scores of longer text spans. For each window, the emotion scores of its constituent words are averaged to obtain the window-level score for that emotion. Teodorescu and Mohammad (2023) showed that emotion arcs constructed with lexicon-based scoring methods, when used with sliding window sizes of 100 instances or more, can mimic the ground-truth emotion arcs with an accuracy of 0.9 or more (higher window sizes also lead to greater accuracies).
Word–emotion scores are obtained from the NRC Emotion Intensity Lexicon, which associates close to 10,000 English words with a real-valued score between 0 and 1 for each dimension. A score of 0 indicates that the word has little to no association with that particular emotion, and a score of 1 indicates a high association.
4.2 Quantifying Emotion Granularity
We compute the emotion granularity metric as the negative of the Spearman correlation between each pair of emotions arcs, for each user. [^7] We choose Spearman correlation as it is rank-based as compared to Pearson correlation which utilizes the raw-values. A high correlation between two arcs indicates that the speaker is consistently and repeatedly expressing the two emotions concurrently; we hypothesize that this is an indicator of a lower ability to differentiate between the two emotions, and therefore a lower emotion granularity. [^8] We choose to use Spearman correlation over ICC-based metrics because the emotion scores that we extract from textual utterances are a relative indicator of the intensity of the emotion, and not an absolute measure. Further, these scores cannot be directly compared across different emotions in terms of absolute intensity (a score of 0.9 for anger may not equate to the same level of anger as a score of 0.9 would for joy) due to differences in how overtly different emotions are expressed via language.
For each person, we average the correlation scores between emotion pairs in the following valence groups to obtain the following measures of emotion granularity (EG):
  • $EG(pos)$: The negative of (i.e., $-1$ times) the average of the correlation scores between each of the pairs of emotions in the positive valence group (joy–trust).
  • $EG(neg)$: The negative of the average of the correlation scores between each of the pairs of emotions in the negative valence group (anger–fear, fear–disgust, etc.).
  • $EG(var)$: The negative of the average of the correlation scores between each of the pairs of emotions in the variable valence group (surprise–anticipation).
  • $EG(overall)$: Overall emotion granularity, measured as the negative of the average of the correlation scores between emotion pairs whose constituents are in the same group. Here, the average is taken across all of the emotion pairs drawn from the positive valence group, the negative valence group, and the variable valence group.
  • $EG(cross)$: Emotion granularity of cross-group emotion pairs. That is, the negative of the average of the correlation scores between emotion pairs whose constituents come from different groups. This measure to some extent quantifies the amount of dialecticism (expressing both positive and negative emotions in a narrow window of time); however, note that EG(cross) also includes emotions that express variable valence (surprise and anticipation), rather than only considering positive–negative valence emotion pairs.
We consider $EG(overall)$ to be the bottom line measure of emotion granularity for a user (analogous to that used in psychology studies). Note that cross-group pairs are not included in this measure.
Dataset, MHC–Control $IC(n)$ $IC(v)$ $EG(pos)$ $EG(neg)$ $EG(var)$ $EG(cross)$ $EG(overall)$
Twitter-STMHD
    ADHD–control lower lower higher lower lower
    Anxiety–control lower lower lower lower lower
    Bipolar–control lower lower lower lower
    MDD–control lower lower lower lower
    OCD–control lower lower lower lower lower
    PPD–control lower lower lower
    PTSD–control lower lower lower lower lower
    Depression–control lower lower lower lower lower
Reddit eRisk
    Depression–control lower lower lower lower lower
Table 2: Emotion Granularity - using Utterance-level Window: The difference in emotion granularity between each MHC group and the control. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference in granularity. E.g., ‘lower’ for ADHD–control under $EG(pos)$ means that the ADHD group has significantly lower positive emotion granularity than the control group.
5 Examining the Link between Emotion Granularity and Mental Health
After preprocessing the posts for each user and computing their emotion granularity scores for the various emotions groups, we performed statistical analyses to determine if there were significant differences between the emotion granularities of each of the MHC groups and the control group.
To test this, we performed an independent t-test between each MHC and the control. However, in order to perform this test the data must meet certain assumptions. These assumptions are explained further in Appendix B. In summary, the data is approximately normally distributed and the groups are independent of each other; however the data does not meet the equal variance (i.e., homoscedasticity) assumption (therefore we specify this when performing the t-test).
We performed independent t-tests to compare emotion granularities between each of the MHCs and the control group for each emotion group using the SciPy library Virtanen et al. (2020). To correct for multiple comparisons (eight tests performed for each MHC per emotion granularity group), we used the Benjamini–Hochberg procedure in the statsmodels library Seabold and Perktold (2010).
5.1 Term Specificity as a Control
Lower emotion granularity occurs when, for a person, the concepts of the relevant emotions are so broad (and non-specific) that their meanings overlap substantially. This work is testing the hypothesis of whether people who have self-disclosed as having an MHC have lower emotion granularity than those that do not. However, another plausible hypothesis is that people in a particular group (e.g., MHC or the control) tend to use more specific words overall. Doing so would imply a higher specificity (i.e, a higher granularity) in their usage of all words, and that the high granularity of emotion words is simply a by-product of their general style of speaking (posting online).
To ensure that the level of word specificity does not differ between MHCs and the control group and act as a confounder for our measure of emotion granularity, we performed a control experiment. We compute the average information content of the noun and verb terms in the posts of users in each group, and use this as a measure of the specificity of their language. The information content of terms has been studied in NLP for applications involving semantic similarity, particularly in information retrieval Resnik (1995); Sánchez et al. (2011).
We use the metric proposed in Resnik (1995), and implemented in the NLTK WordNet library, [^9] https://github.com/nltk/wordnet which combines information about the depth of the term in the WordNet tree hierarchy and its frequency of occurrence in a large corpus (here, the Brown corpus) to compute an information content score Miller (1995). We lemmatize each word using the inbuilt WordNet Morphy library, and consider the most frequent synset to obtain the information content score. We then compute the following measures of term specificity:
  • $IC(n)$: The information content score for all nouns is averaged across all posts of each individual in each group
  • $IC(v)$: The information content score for all verbs is averaged across all posts of each individual in each group. Statistical tests for significant differences are similarly performed as described above (Section 5).
  • 6 Results
    In Table 2 we report the statistical results from the pairwise comparisons between each MHC and the control group, for the control experiment on general term specificity as well as emotion granularity, when scores are computed at the utterance-level. All statistically significant differences between an MHC and the control group are described as either ‘higher’ or ‘lower’, and a dash (–) for no statistical difference. A ‘lower’ value in a cell indicates that the MHC (rows) has lower emotion granularity (or lower term specificity) than the control group, i.e., higher correlation between emotion pairs in that group (columns); ‘higher’ indicates the MHC has higher emotion granularity (or higher term specificity) than the control group (i.e., lower correlation between emotion pairs in that group). Below we summarize the results for each column.
    IC(n) and IC(v): We do not see any significant differences in the information content of noun and verb terms ($IC(n)$ and $IC(v)$) between MHCs and the control group. This indicates that no group tends to use more specific or less specific language in general when posting on these platforms. Details on the statistical results are shown in Appendix E.
    EG(pos): All MHCs except for PPD had significantly lower positive emotion granularity than the control group (which had similar granularity compared to the control group). That is, these MHCs (ADHD, Anxiety, etc.) consistently expressed multiple positive emotions concurrently, more so than the control group.
    EG(neg): All MHCs had significantly lower negative emotion granularity than the control group, in both datasets. Thus these MHCs were generally not differentiating between the negative emotions of anger, disgust, fear, and sadness, as well as compared to the control group.
    EG(var): All MHCs except ADHD had significantly lower variable emotion granularity than the control group (i.e., these MHCs generally differentiated between surprise and anticipation less than the control group). On the other hand, ADHD had significantly higher variable emotion granularity than the control group (i.e., the ADHD group was better at differentiating between surprise and anticipation).
    EG(overall): All MHCs had significantly lower emotion granularity for emotion categories that express the same valence (the mixed valence emotions of surprise and anticipation are also included here). These MHCs are therefore expressing multiple close emotions frequently with one another – more so than the control group.
    EG(cross): ADHD, Anxiety, OCD, PTSD, and Depression (in both datasets) have significantly lower granularity between emotion pairs that come from different valence groups. This indicates that positive and negative emotions are expressed together more frequently in these MHCs compared to the control, as well as emotions like joy (positive valence) and surprise (variable valence).
    Discussion: These results demonstrate that our measures of emotion granularity from text are consistently lower for users in the MHC groups compared to the control. The term specificity results also tell us that it is the specificity of emotion word usage in particular that is differentiating MHCs from the control group.
    Aligning with self-report studies in psychology, the emotion granularity between negative-valence emotions is lower for all MHCs with utterance-level operationalizations. Positive emotion granularity is also correlated with many of the MHCs (except PPD). In general, the granularity of emotional expression between within-group emotion pairs is lower for all MHC groups compared to the control in both datasets. This is in line with both the theoretical and conceptual links established in the psychology literature on emotion granularity and mental health: the ability to better differentiate between emotion concepts that are close to one another leads to more adaptive health behaviour.
    While emotion pairs from differently-valenced emotion groups are not usually operationalized in affective science experiments, we find that this measure is also significantly lower in many MHCs. Further investigations into what the concurrent expression of positive and negative emotions means, for emotion granularity and emotion dynamics in general, are interesting research directions.
    6.1 Additional Window Sizes
    We examine how the measure of emotion granularity between MHCs and the control changes when we compute emotion scores with larger window sizes (multiple tweets/posts are included in computing the emotion score at a single time point).
    Many of the utterance-level outcomes are replicated for negative, positive, and overall emotion granularity with window size 100. Some measures are no longer significantly different between certain MHCs and the control (e.g., $EG(overall)$ for PPD). We also find that $EG(cross)$ is higher for certain MHCs (Anxiety, PPD, PTSD, Depression in Twitter-STMHD) when compared to the control, i.e, users in the control group are expressing negative and positive emotions together more frequently than those in the MHCs. When the window size is increased to 500 words, this reversal of trends for cross-valence emotions persists.
    With larger window sizes, we end up capturing emotions expressed by the individual over longer time spans (tweets posted over the span of several hours or days), rather than co-endorsement at the same time. We hypothesize that these effects of dialecticism, where the Control group has a higher co-occurrence of cross-valence group emotions, are capturing the extent to which users balance negative emotions with positive emotions (and vice versa), and how people bounce back from negative events, find the silver linings, etc. The consistent effects with 100 and 500-word windows, and for several MHCs, makes this a promising area for future work.
    All emotion granularity measures with window sized 100 and 500 are reported in Tables 3 and 4 in Appendix C.
    6.2 Individual Emotion Pairs
    In order to understand which emotion pairs are expressed together more frequently (resulting in lower emotion granularity), we performed the same significance tests as before between MHCs and the control for correlation scores between all individual emotion pairs. We found that:
    • All eight MHCs in the Twitter-STMHD dataset had a lower granularity (a higher correlation) for anger–disgust and anger–sadness in the negative valence group.
    • Among cross-group emotion pairs, all eight MHCs had a lower emotion granularity (higher correlation) between anticipation–fear and anticipation–sadness.
    • Certain MHCs had a higher emotion granularity (lower correlation) between the following cross-group emotions: anger–joy co-occurs more in the control than in Bipolar, Depression, MDD, PPD, PTSD; however, it co-occurs less in control compared to ADHD.
    • Similarly, sadness–surprise co-occur more in the control when compared to Anxiety, Bipolar, Depression, and PPD; but less when compared to ADHD and PTSD.
    Detailed results for each of the emotion pairs and all MHCs are in Appendix D, Table 5.
    Discussion: While lower granularity among certain emotion pairs consistently function as indicators of all MHCs, we also see many instances where MHCs have a higher granularity between the emotions when compared to the control. These findings are of interest to researchers studying the links between how emotions are expressed in text, and how they are vary with with different MHCs.
    7 Conclusion
    In this work, we operationalized for the first time a computational measure of emotion granularity that can be derived from the textual utterances of individuals. We applied this measure to two social media datasets of posts from individuals who have self-disclosed as having an MHC. Our findings showed that our measure of emotion granularity is significantly lower for all the MHC groups when compared to a control group at an aggregate-level, for emotion pairs that express similar valence. Specifically, we found significantly lower values of negative and overall emotion granularity for all eight MHCs represented in our dataset. Similarly, we found all MHCs except for PPD had lower positive emotion granularity than compared to the control group. Our work makes an important contribution towards deriving aggregate-level indicators of emotional health from the large amounts of utterance data available on social media platforms.
    Limitations
    Our work uses the social media utterances of individuals to derive measures of emotional expression that, at an aggregate level, are found to correlate with multiple mental health conditions. While we use datasets that were compiled by other researchers in the field, we stress that they may not be representative of the general population. Our methods therefore cannot be directly applied to make inferences on other datasets without a careful experimental validation first. The datasets we study rely on self-disclosures made on social media platforms; it is possible that users report only one such MHC but are diagnosed with others, or that they misrepresent their diagnoses. Further, the users in the control groups may include those who have chosen to simply not self-disclose on these platforms. This can occur due to many reasons, like social desirability Latkin et al. (2017b) or impression management Tedeschi (2013)). Nonetheless, since we draw inferences at an aggregate level, the methods used can overcome some amount of noisy data.
    The set of emotions that we have considered in our measurement of emotion granularity are also limited to those for which we can computationally obtain text-derived emotion scores. These eight emotions do not represent the wide range of emotion concepts that exist and are experienced and expressed by us with language, and future research can attempt to expand our operationalization to more emotion concepts. It should be noted though, that past psychology studies on emotion granularity have also tended to explore small sets of emotions, largely because it is cumbersome to ask users about how they feel for a large set of emotions.
    The emotion lexicons that we use are some of the largest that exist with wide coverage and large number of annotators (thousands of people as opposed to just a handful). However, no lexicon can cover the full range of linguistic and cultural diversity in emotion expression. The lexicons are largely restricted to words that are most commonly used in Standard American English and the capture emotion associations as judged by American native speakers of English. See Mohammad (2020) for a discussion of the limitations and best-practises in the use of emotion lexicons.
    Ethics Statement
    Our approach, as with all data-driven models of determining indicators of mental health, should be considered as aggregate-level indicators, rather than biomarkers for individuals Guntuku et al. (2017). We do not attempt to predict the presence of MHCs for individual users at any stage of the process. These measures should also not be taken as standalone indicators of mental health or mental wellness, even at the population level, but rather as an additional metric that can be used in conjunction with other population-level markers, and with the expertise of clinicians, psychologists, and public health experts.
    Individuals vary considerably in how, and how well, they express their internal emotional states using language. Our method of assessing the emotional states of users based on their utterances may miss several linguistic cues of emotion expression, and may not account for individual variation or the extent to which these emotions are expressed on social media. The emotionality of one’s language may also be conveying information about the emotions of the speaker, the listener, or something or someone else mentioned in the utterances. See further discussions of ethical considerations when using computational methods for affective science in Mohammad (2020, 2022).
    Appendix A Twitter-STMHD Dataset
    Suhavi et al. (2022) created a regular expression pattern to identify posts which contained a self-disclosure of a diagnosis and the diagnosis name (using a lexicon of common synonyms, abbreviations, etc.) such as ‘diagnosed with X’. They collected a large set of tweets using the regex. This resulted in a preliminary dataset of users with potential MHC diagnoses. To handle false positives (e.g., ‘my family member has been diagnosed with X’, or ‘I was not diagnosed with X’), the dataset was split into two non-overlapping parts, one of which was manually annotated, and the other using an updated and high-precision regex. In the part that was annotated by hand, each tweet was annotated by two members of the team. A user was only included in the dataset if both annotations were positive as self-disclosing for a particular class. A licensed clinical psychologist found the 500-tweet sample to be 99.2% accurate. The manual annotations were used to refine the regular expressions and diagnosis name lexicon. This updated search pattern was applied to the other dataset split. To verify the quality of the updated regex, the authors applied it to the manually annotated dataset split. When considering the manual annotations as correct, the regex was found to be 94% accurate.
    Appendix B Statistical Assumptions
    Below we describe in more depth the requirements for performing an independent t-test, which was done in our analyses.
    • The dependent variable must be measured using a continuous scale: emotion granularity is measured as the average of Spearman correlation between emotion arcs in the group, resulting in continuous values.
    • The independent variable must have two categorical and independent groups: Our independent variable is diagnosis, which is either an MHC or the control group.
    • Independence of observations: Since the text stream of utterances come from different people, we can assume these are independent observations. [^10] In reality individuals are largely influenced by one another as we see interact and engage with content from various communities, which can influence our emotional state and therefore utterances. However, for the purposes of our experiments since the utterances come from different people we can assume they are independent.
    • Approximately normally distributed dependent variable for each group of independent variable: Given the large number of people and number of utterances per person in our dataset, we can assume that the means of the data for each group is approximately normally distributed according to the law of large numbers. Further, the t-test is robust to violations of normality.
    • Homogeneity of variance: We performed Levene’s test for homogeneity of variance to verify whether this assumption is met. Our data did not meet this assumption, therefore we performed t-tests with the unequal variance setting as True in SciPy.
    Appendix C Emotion Granularity: Various Window Sizes
    We report the results of the statistical analyses of emotion granularity when emotion arcs were generated using two other window sizes: 100 (Table 3) and 500 (Table 4). We find that largely the results do not change, however there are some differences in the scenario when the dataset was smaller (e.g., eRisk dataset or MHC such as MDD). In such cases, when the window size is increased, it is possible that several emotional experiences occurred, resulting in a weaker signal of emotion granularity.
    Dataset, MHC–Control $IC(n)$ $IC(v)$ $EG(pos)$ $EG(neg)$ $EG(var)$ $EG(cross)$ $EG(overall)$
    Twitter-STMHD
        ADHD–control lower lower higher lower lower
        Anxiety–control lower lower lower higher lower
        Bipolar–control lower lower lower lower
        MDD–control lower lower lower
        OCD–control lower lower lower
        PPD–control lower higher
        PTSD–control lower higher lower
        Depression–control lower lower lower higher lower
    Reddit eRisk
        Depression–control lower lower lower
    Table 3: Emotion Granularity - using Window 100: The difference in emotion granularity between each MHC group and the control. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference in granularity.
    Dataset, MHC–Control $IC(n)$ $IC(v)$ $EG(pos)$ $EG(neg)$ $EG(var)$ $EG(cross)$ $EG(overall)$
    Twitter-STMHD
        ADHD–control lower lower lower lower
        Anxiety–control lower lower lower higher lower
        Bipolar–control lower lower lower lower
        MDD–control lower
        OCD–control lower
        PPD–control higher
        PTSD–control lower higher lower
        Depression–control lower lower lower higher lower
    Reddit eRisk
        Depression–control lower
    Table 4: Emotion Granularity - using Window 500: The difference in emotion granularity between each MHC group and the control. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference in granularity.
    Appendix D Emotion Granularity: Emotion Pairs
    In Table 5 we report the pairwise emotion granularity results when testing for significant differences between MHCs and the control group.
    Emotion Pair, MHC-Control ADHD Anxiety Bipolar Depression MDD OCD PPD PTSD
    anger–anticipation lower lower lower lower lower lower
    anger–disgust lower lower lower lower lower lower lower lower
    anger–fear lower lower lower lower lower
    anger–joy lower higher higher higher higher higher
    anger–sadness lower lower lower lower lower lower lower lower
    anger–surprise lower higher lower higher lower
    anger–trust lower lower lower lower lower
    anticipation–disgust lower lower lower lower lower lower
    anticipation–fear lower lower lower lower lower lower lower lower
    anticipation–joy lower lower higher
    anticipation–sadness lower lower lower lower lower lower lower lower
    anticipation–surprise higher lower lower lower lower lower lower lower
    anticipation–trust lower lower lower lower lower
    disgust–fear lower lower lower lower lower lower lower
    disgust–joy lower higher higher higher
    disgust–sadness lower lower lower lower lower lower
    disgust–surprise lower lower lower lower higher lower
    disgust–trust lower lower lower lower lower
    fear–joy lower
    fear–sadness lower lower lower lower lower lower lower lower
    fear–surprise lower higher higher higher
    fear–trust lower lower lower lower lower lower
    joy–sadness lower higher
    joy–surprise higher lower lower
    joy–trust lower lower lower lower lower lower lower
    sadness–surprise lower higher higher higher higher lower
    sadness–trust lower lower lower lower lower
    surprise–trust lower lower lower lower lower lower lower
    Table 5: Emotion Granularity - Utterance-Level: The difference in emotion granularity between each emotion pair, for each MHC group and the control in the Twitter-STMHD dataset. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference.
    Appendix E Term Specificity Results
    Table 6 shows the results of the term specificity experiments described in Section 5.1 measuring information content. For both nouns and verbs, none of the diagnoses had significantly different term specificity levels compared to the control group in both the Twitter-STMHD and eRisk datasets. This verifies that the significant differences between the MHCs and the control group for emotion granularity is not due to varying word specificity levels in these groups.
    df T-Statistic P-value
    POS Dataset MHC–Control
    Noun Twitter-STMHD ADHD–control 2368.64 -1.94 0.144
    Anxiety–control 2233.36 -0.58 0.718
    Bipolar–control 1726.26 -2.40 0.131
    MDD–control 226.83 -0.52 0.718
    OCD–control 1817.93 1.93 0.144
    PPD–control 178.33 -0.49 0.718
    PTSD–control 2237.82 -0.54 0.718
    Depression–control 2245.64 -0.27 0.787
    Reddit eRisk Depression–control 128.95 0.98 0.330
    Verb Twitter-STMHD ADHD–control 2248.45 -0.73 0.530
    Anxiety–control 2235.85 1.12 0.420
    Bipolar–control 1852.44 -2.10 0.096
    MDD–control 213.0 1.36 0.354
    OCD–control 1645.17 2.28 0.091
    PPD–control 169.49 0.98 0.438
    PTSD–control 2351.12 2.53 0.091
    Depression–control 2274.59 0.54 0.589
    Reddit eRisk Depression–control 110.40 1.59 0.116
    Table 6: Information Content: The degrees of freedom, t-statistic and p-value for the word specificity experiments described in Section 5.1.
    Appendix F Emotion Correlations
    Table 7 shows the Spearman correlation between emotion arcs for all pairs of emotions for the control group. These results indicate that as baselines largely emotions in the same group (e.g., positive, negative, mixed, overall) co-occur more often than emotions across groups.
    Table 8 shows the group-averaged Spearman correlations for emotion pairs in the positive, negative, mixed valence groups, and the within-group and cross-group averages, for the MHCs and Control groups in both datasets.
    emo2 anticipation disgust fear joy sadness surprise trust
    emo1
    anger 0.020 0.593 0.536 -0.015 0.447 0.113 0.003
    anticipation 0.009 0.040 0.432 0.022 0.251 0.404
    disgust 0.429 -0.014 0.434 0.061 0.001
    fear -0.010 0.567 0.138 0.004
    joy -0.009 0.221 0.467
    sadness 0.100 -0.001
    surprise 0.145
    Table 7: Emotion–Emotion Spearman Correlation: Spearman correlation values between pairs of utterance-level emotion arcs for the control users in the Twitter-STMHD dataset.
    Dataset, MHC $EG(pos)$ $EG(neg)$ $EG(var)$ $EG(cross)$ $EG(overall)$
    Twitter-STMHD
        Control 0.467 0.501 0.251 0.083 0.465
        ADHD 0.495 0.524 0.243 0.096 0.486
        Anxiety 0.505 0.522 0.274 0.089 0.489
        Bipolar 0.491 0.524 0.273 0.084 0.488
        Depression 0.505 0.520 0.272 0.087 0.487
        MDD 0.508 0.518 0.267 0.085 0.486
        OCD 0.497 0.523 0.260 0.090 0.487
        PPD 0.473 0.511 0.269 0.081 0.476
        PTSD 0.485 0.521 0.259 0.090 0.484
    Reddit eRisk
       Control 0.475 0.537 0.271 0.168 0.496
       Depression 0.521 0.568 0.319 0.184 0.531
    Table 8: Emotion Groups Spearman Correlation: Spearman correlation values between utterance-level emotion arcs of emotion pairs in the positive, negative, and variable valence groups, and within-group (overall), and cross-group averages, for the Twitter-STMHD and Reddit eRisk datasets. Emotion granularity is defined as the negative of these correlations (i.e, higher correlations imply a lower granularity).